bill533:

A square pyramid has a volume of 108 cubic feet and a height of 4 feet. What is the length of each side of the base of the pyramid? A. 4 feet B. 9 feet C. 18 feet D. 27 feet E. 81 feet

AngeI:

V=(a^2)(h/3)

AngeI:

108=(a^2)(4/3) 108=(a^2)1.333

AngeI:

Then i believe you just have to divide 1.333 from both sides then find a

bill533:

D.

AngeI:

Nooo Divide both sides by 1.33 and you get 81=a^2

AngeI:

So a equal 9